At January 1, 2020, Betty DeRose, Inc. reported work in process totaling $18,670. This balance was made up of three jobs haviCalculate Betty's:

A) Work In Process Inventory Balance at December 31, 2020

B) Finished goods inventory balance at December 31, 2020.

C) gross profit reported in Betty DeRose's income statement for 2020.
Answer #1

Job Cost Sheet

Job 171 Job 236 Job 412
Beginning balance 4390 6370 7910
Direct material 6700 4100 3900
Direct labor 5000 2800 3000
Overhead 5695 3485 3315
Total 21785 16755 18125
Unit 100
Unit price 217.85

Applied overhead = 5695+3485+3315 = 12495

Actual overhead = 10000

Over applied overhead = 2495

Adjusted cost of goods sold = (40*217.85)-2495 = 6219

Gross profit = Sales-Cost of goods sold = 12000-6219 = 5781

a) Work in process inventory balance at December 31,2020 = 16755

b) Finished goods inventory = 60*217.85+18125 = 31196

c) Gross profit = 12000-6219 = 5781